Lung Cancer & Smoking Case Study A causal relationship between cigarette smoking and lung cancer was first suspected in the 1920s on the basis of clinical observations. To test this apparent association, numerous epidemiologic studies were undertaken between 1930 and 1960. Two studies were conducted by Richard Doll and Austin Bradford Hill in Great Britain The first was a case-control study begun in 1947 comparing the smoking habits of lung cancer patients with the smoking habits of other patients. The second was a cohort study begun in 1951 recording causes of death among British physicians in relation to smoking habits. This case study deals first with the case- control study, then with the cohort study. Data for the case-control study were obtained from hospitalized patients in London and vicinity over a 4-year period (April 1948 - February 1952). Initially, 20 hospitals, and later more, were asked to notify the investigators of all patients admitted with a new diagnosis of lung cancer. These patients were then interviewed concerning smoking habits, as were controls selected from patients with other disorders (primarily non-malignant) who were hospitalized in the same hospitals at the same time Data for the cohort study were obtained from the population of all physicians listed in the British Medical Register who resided in England and Wales as of October 1951. Information about present and past smoking habits was obtained by questionnaire. Information about lung cancer came from death certificates and other mortality data recorded during ensuing years Over 1,700 patients with lung cancer, all under age 75, were eligible for the case-control study. About 10% of these persons were not interviewed because of death, discharge, severity of illness, or inability to speak English. An additional group of patients were interviewed but later excluded when initial lung cancer diagnosis proved mistaken. The final study group included 1,530 casesTable 2 shows the frequency distribution of male cases and controls by average number of cigarettes moked per day Table 2. Most recent amount of cigarettes smoked daily before onset of the present illness, lung cancer cases and matched controls with other diseases, Great Britain, 1948-1952 Daily number of cigarettes # Cases # Controls Odds Ratio 68 785 482 195 1,462 1,530 referent 1-14 15-24 25+ All smokers Total 622 498 402 1,522 1,530 Question 14 Compute the odds ratio by category of daily cigarette consumption, comparing each smoking category to nonsmokersQUESTION 15: WHAT DO THESE RESULTS TELL YOU?

QUESTION 16: WHAT ARE THE OTHER POSSIBLE EXPLANATIONS FOR THE APPARENT ASSOCIATION? INCLUDE AND EXPLAIN 3 SPECIFIC BIAS THAT PERTAIN TO THIS STUDY.
Answer #1

14) smoking category to nonsmokers: 1-14 daily cigarette: 622*68/785*8=7.0, 15 -24 daily cigrattes : 498*68/482*8=8.7, 25 + cigarettes :402*68/195*8=17.5, all smoker=1522*68/1462*8=8.8. 15)As the number of cigarettes increases we can see an increase in the odd of lung cancer and we can see the association between smoking the number of daily cigarettes and lung cancer , 16) Bias such as selection and observational could investigator or simply chance error could also affect the results. and other possible explanation for the apparent association between smoking and lung cancer may include information or selection bias confounding , an error on the part of the investigator or simply chance although the data indicated that chance is not likely.